hello, can someone please explain how to find the derivative of q=100/P ?

and what is the derivative itself

thank you
Answer #1

Solution: The derivative of a function q = f(p) of a variable p is a measuring the rate at which the value q of the function changes in regard to the change of the variable p. It is known as the derivative of f with respect to p.

 

Derivative of q=100/P

q = 100P^-1

Differentiate using the Power Rule,Pn-1 1 =-1*P 1000 P2